### v1

@@ 4 Results
In contrast, the high-renewable scenarios, including both zero-emissions and socially-optimal scenarios with a $190/tCO, have starkly greater variability, with 28 to 48 percent of delivered MWh having a zero or near-zero marginal cost ( $1/MWh). The large share of “free" energy occurs because solar and wind capacities are optimally built out to help serve demand on days with low to medium wind and sunlight and high demand. These times are seasonal, tending to occur in late fall and winter. As a result, there are often surpluses during other seasons, especially spring and summer, when wind and sunlight are more abundant, leading to many stretches of time with zero marginal cost. During these times, renewables are curtailed because there is more energy in batteries than could be depleted before being fully charged again with excess wind and solar, and electrolysis capacity is fully utilized.
@@
In zero-emissions scenarios, the spread of marginal cost above near-zero is smooth and broad, with about 25 percent of marginal costs spread evenly between $10 and $100 per MWh, and the marginal cost of most other energy spread evenly between $100 and $200 per MWh. Extreme marginal costs above $1000/MWh are rare. The main effect of transmission is to slightly increase the frequency of hours in the $30 to $80 and $100 and $200 range. Again, the influence of transmission is modest. In the socially optimal scenarios, marginal costs above near-zero are more concentrated in the $75 to $200 range. Like the other cases, the influence of transmission on the distribution of marginal costs is minimal.
@@
The extreme variability of marginal cost in high-renewable scenarios, especially the high frequency of zero-marginal-cost energy, suggests that flexible demand could be especially valuable in these systems, as indicated by (Imelda, Fripp and Roberts, 2022).
@@
Transmission expansion increases economic rents for some resource owners while reducing them for others (Figure 6). If resource owners are paid a price equal to the the incremental cost of power in each region and hour, then all variable and fixed costs are recovered in long-run equilibrium, which is what our model assumes. In addition, some owners of scarce resources glean rent above their total costs. The potential rents we can identify in this study pertain to heterogeneous solar and wind resources and potentially constrained transmission resources. Areas with unusually large or well-timed amounts of solar radiation or winds earn rent (above competitive profit) because they provide more value than marginal resources. Our model indicates that these rents tend to be highest in the zero-emissions scenarios.
@@
Combining rent and cost changes, which mostly benefit consumers, all regions gain from transmission expansion. However, since some resource owners lose rents relative to the no-expansion case, diverging stakeholder interests could hamper efforts to expand transmission. Figure 6 illustrates these divergent interests for solar and wind resources, which are most prevalent and have greater rent than in the zero-emissions scenarios. The graph shows gains and losses in solar, wind, nuclear, and transmission rent. The coastal regions and South tend to lose solar and wind rent with transmission expansion, while Arizona, Ohio, and Michigan gain solar rent, and the Rocky Mountains, Oklahoma, and upper Great Plains gain wind rent. Nuclear plants mostly lose rent, as sunk costs from existing plants earn less revenue when transmission is optimized, enhancing competition from renewables.
@@
Hourly operational detail in ERCOT shows how transmission expansion can aid resiliency during the most constrained times (Figure 7). The model optimizes hourly operations and interregional flows in each region, revealing how the system achieves balance. Here we consider a costly day in the most isolated region, TRE, the Eastern part of ERCOT, which has most of the interconnect’s generation and demand. In two of the four scenarios (LO, ZE), the day depicted (December 13th, 2050, based on actual weather data in 2012), is the most expensive in the region. This is a cold day in early winter near the solstice, when solar resources would generally be more limited, with unusually little wind.
@@
Compared with the least-cost scenarios, decarbonization is achieved via substantially greater capacities of solar and wind paired with significant use of batteries and hydrogen, which replaces natural gas and some old-vintage coal. The deficit from renewable energy is mostly counterbalanced with natural gas in the least-cost scenarios and with imports, battery storage, and hydrogen in the zero-emissions scenarios. Notably, even the least-direct-cost scenarios use batteries to help serve the net peak at 7pm, when wind and solar power fall to near-zero.
@@
The effects of optimizing transmission are subtle and difficult to discern in the least-direct-cost scenarios. In these scenarios, some power is imported from the Western part of ERCOT (TRE_WEST) at 1pm and 3pm of December 13th, 2050 when transmission is fixed (panel a), but less is imported under optimal transmission, because, in this case, more natural gas capacity is built so the region can export more on other days. The generation mix is largely the same when transmission is fully optimized (panel b), except there is slightly more gas generation and less battery use during the evening peak. The peak hour marginal cost falls from about $20,000 to about $16,000 when optimizing transmission.
@@
The zero-emissions scenarios (c and d) show a somewhat greater influence from optimizing transmission. Generation mixes are similar regardless of transmission, but there is noticeably more battery use and slightly less wind in the scenario with existing transmission. And while imports help, especially in the optimal transmission scenario, it is interesting that imports occur midday to help replenish batteries, and much less during the expensive evening peak. The peak marginal cost of about $22,000 per MWh falls to about $7,000 with expanded transmission. And where the nighttime hours after the peak stay elevated at $5,000 per MWh with existing transmission, marginal cost falls back to very low levels when transmission is optimally expanded.
@@
Although we do not consider demand response or demand-side flexibility in this study, this difficult day in ERCOT suggests it could be valuable. In all scenarios, hours up to 5pm are very low cost or free. Adjustments in EV charging, thermal storage or pre-heating in buildings, or even modest thermostat adjustments during the critical peak, would reduce required hydrogen or battery capacity, saving substantial costs. Even existing demand responses by some industrial customers could make a difference. Such demand responses would likely reduce transmission value, too.
@@ 5 Discussion
More than 2 TW of solar, wind, and storage capacity was waiting for interconnection to the U.S. grid at the end of 2022, a staggering quantity equal to multiple times existing capacity. The U.S. electricity sector would take a healthy step toward decarbonization if all of it were built, connected to the grid, and managed efficiently. Still, the wait times are substantial, and many proposed projects are never completed (Rand et al., 2023). Full decarbonization will require even more growth of renewable energy capacity by 2050, especially with higher demand from electric vehicles and electrification of heat and industry. Many kinds of local transmission upgrades are critical to connect these resources, including spur lines to the locations of new generation facilities, and grid-enhancing technologies that can help to maximize the use of existing transmission lines within and between regions. The Federal Energy Regulatory Commission is considering the nature and extent of transmission upgrades needed, and how best to manage a considerable queue of clean energy resources.
This study does not focus on these immediate transmission challenges of grid interconnection and optimal power flow (which considers exact existing lines, their impedances, currents, voltage limits, and real and reactive power flows). However, it does account at least approximately for the cost of spur lines and transmission upgrades to connect new solar and wind installations to the grid. Rather, considered the costs and benefits of transmission expansion and large-scale energy flows between large regions, which could allow more power to be traded over greater distances but could face considerable social, political, and environmental challenges. This research question and scope can be adequately addressed using capacity expansion models as we do in this work.
@@
While the optimal expansion of interregional transmission would require more than a three-fold increase relative to what presently exists, we find that the net benefits of this expansion are small relative to the overall cost of a decarbonized system. Alternatives to expanded transmission are multifaceted and include greater renewable generation capacity, battery storage, and hydrogen electrolysis paired with fuel cell power generation. While expanded transmission might improve resiliency in some locations, the overall distribution of marginal costs (i.e., competitive wholesale prices) is similar regardless of interregional transmission expansion.
@@
One caveat is that some kinds of transmission might be expanded at lower cost using various grid enhancing strategies111https://www.energy.gov/oe/grid-enhancing-technologies-improve-existing-power-lines to improve power flow, or by employing new high-capacity conductors in existing right-of-ways (Chojkiewicz et al., 2024), which could halve the cost of expanding existing lines. It is also conceivable that storage costs could fall more than projected. Such technologies could only serve to expand the portfolio of options available to integrate clean energy affordably.
@@
Another caveat to these findings is that they generally assume economic efficiency or perfectly competitive outcomes. During congested times, transmission and generation capacity owners may possess considerable market power, and be tempted to withhold generation or transmission to raise wholesale prices and profits (Borenstein, Bushnell and Wolak, 2002). Greater transmission capacity could, therefore, have the added benefit of increasing competition. At the same time, those earning substantial scarcity and/or market-power rents may resist expansion efforts that would reduce such rents. Transmission expansion may need to be expanded much more than otherwise optimal if the goal is to attenuate market power.
@@
Perhaps a greater limitation of this study is that it does not account for demand-side adjustments that could substitute for transmission, storage, and generation. Such responses could be achieved through strategic timing of electric vehicle charging, vehicle-to-grid technologies, improved building insulation, thermal storage, smart control of building HVAC and water heating systems, thermostat adjustments during extreme weather events, and industrial demand shifting, some of which already exist. Given the extraordinary growth in marginal cost variability in the decarbonization scenarios, it seems likely that utilities, industry, communities, and opportunistic intermediaries will find ways to exploit growing wholesale price differentials. If they do, interregional transmission growth could be worth less than we estimate. Future research could consider how optimal transmission upgrades and expansion interact with demand response and distributed resources from local to interregional scales. A robust demand-side market that is responsive to real-time marginal costs would have the added benefit of reducing potential market power during congested times.
@@ 6 Methods
Limitations
Three main aspects are important to discuss in the context of transmission line expansion. First, in this study, we model transmission capacity in an aggregated manner to represent important energy flows between load zones. Our approach models the transfer capability of the electrical network but does not directly represent the physics of AC optimal power flow. Hence, our findings should not be taken as a claim that there is no need to worry about transmission expansion within individual regions in 2050. Our work sheds light on what are the most important electricity transfers between regions and interconnects to achieve zero emissions in 2050. Our findings speak to the lack of need and minimal economic impact of deploying “energy corridors” between interconnects as opposed to speaking about the value of specific transmission lines, for which power flow modeling would be required with an accurate model of the transmission network (e.g., California should be modeled with its approximately 9,000 buses). Second, our study, while high-resolution, considers only a single weather year and coincident demand. We do not account for reserves or for extreme weather events that do not occur in a typical year. Appropriate account of reserves and extreme events in reliable high-renewable environments requires the compilation of many years of high-resolution weather. It would also be important to include a careful account of demand response in such a study.
@@
Our analysis uses Switch (Fripp, 2012; Johnston et al., 2019), which is an open source capacity expansion model that minimizes the net present value (NPV) of costs for an electricity grid across all investment periods and timepoints. It optimizes both the investment and operational costs subject to constraints such as emissions limits or prices, or other policies, such as Renewable Energy Portfolio Standards (RPS).
@@
One of the most difficult parts of running electricity capacity expansion models is assembling all the data. PowerGenome (Schivley, 2020) serves as a platform to generate inputs files for power system optimization models—including Switch, GenX, Temoa and USENSYS. The source data comes from a number of different resources, including EIA, NREL, and EPA.
@@
Existing generating units are from the latest version of form EIA-860 (U.S. Energy Information Administration, 2021), 2021 supplemented with 860m from June 2023. The cost and heat rates of new-build resources are provided by NREL’s Annual Technology Baseline (ATB) 2022 (National Renewable Energy Laboratory, 2022). Renewable resources are generally location-specific, with unique generation profiles and interconnection costs. Rather than representing all potential new-build renewable resources as individual sites, PowerGenome lets the user specify how much capacity of each resource type should be available for consideration in a model region and how many clusters the resource should be represented by. We use 942 clusters of candidate renewable energy projects, corresponding to approximately 21 TW of potential capacity. High-resolution weather data provided by Vibrant Clean Energy for 2012 (the year upon which our demand data is based) is used to construct hourly variable capacity factors for each cluster.
@@
Existing transmission capacities between IPM regions are from EPA. Model regions in our study can consist of one or more individual IPM regions. When two or more IPM regions are combined to make a Switch model region, the transmission capacity between individual IPM regions is also combined. Transmission lines in this study are defined as connections between major metros in each region, with additional backbone networks connecting major metros within a region (if there is more than one). Costs and line loss use transmission line segments created with a least cost path method described in Appendix F of (Patankar et al., 2023).
@@
Hourly demand starts with NREL EFS profiles (National Renewable Energy Laboratory, 2019). Stock values (historical) and hourly demand profiles of electrified end-use technologies like transportation, water heating, and space heating/cooling (derived from NREL EFS data) are subtracted from the 2019 EFS profiles. The remaining demand is inflated using sector-specific growth rates from EIA AEO 2022 (Nalley and LaRose, 2022). The future hourly demand from electrified end-use technologies are added back in using future stock values (from the REPEAT (Rapid Energy Policy Evaluation and Analysis Toolkit, 2022) scenario IRA_MID) and hourly demand profiles derived from EFS.
@@
Supplemental Analysis
Here we describe mathematical structure of Switch in more detail and present additional results, with an emphasis on comparing the least-direct-cost (LX) scenarios with the socially optimal (SX) scenarios that minimize costs subject to a carbon emissions price of $190 per ton rather than forcing zero emissions. This price leads to 90 to 95% emissions reductions relative to 2022 levels in most cases, even with a 75% growth in electricity production. We refer to these scenarios as socially optimal because they equate the marginal cost of abatement to the Environmental Protection Agency’s marginal social cost of emissions (i.e., the estimated marginal benefit of abatement).
@@
Here we provide a brief overview of the mathematical formulation of Switch 2.0, the model used for our research. More complete documentation of the software can be found in (Johnston et al., 2019). Switch 2.0 has a modular architecture that reflects the modularity of actual power systems. Most power system operators follow rules that maintain an adequate supply of power, and most individual devices are not concerned with the operation of other devices. Similarly, core modules in Switch define spatially and temporally resolved balancing constraints for energy and reserves, and an overall social cost. Separate modules represent components such as generators, batteries or transmission links. These modules interact with the overall optimization model by adding terms to the shared energy and reserve balances and the overall cost expression. Switch 2.0 supports co-optimization of multiple investment periods, but we have omitted those definitions here, since we use a single stage for this study (2041-2050). We have likewise omitted details on spinning reserves and unit commitment, which were not used for this study.
@@
Additional terms define behavior of storage facilities (charging, state of charge, round-trip losses), fuel consumption (full load heat rate times power production) and hydrogen facilities (production and storage of hydrogen and conversion back to electricity). For complete details on Switch’s mathematical formulation, see the Supplementary Material of (Johnston et al., 2019) at https://ars.els-cdn.com/content/image/1-s2.0-S2352711018301547-mmc1.pdf.
@@
Least-direct-cost | Zero-emissions | Socially-optimal | |||||||
LE | LI | LO | ZE | ZI | ZO | SE | SI | SO | |
O&M | 4.2 | 4.2 | 4.2 | 1.8 | 1.4 | 1.4 | 5.5 | 4.9 | 4.6 |
Fuel | 35.3 | 34.9 | 34.5 | 4.8 | 3.6 | 3.7 | 15.9 | 13.8 | 13.3 |
Storage | 1.4 | 1.3 | 1.3 | 15.8 | 14.2 | 13.0 | 7.4 | 6.37 | 6.0 |
Generation | 36.3 | 35.6 | 35.4 | 100.1 | 92.9 | 91.9 | 77.6 | 74.6 | 73.9 |
Transmission | 3.1 | 4.0 | 4.3 | 3.1 | 9.1 | 9.8 | 3.1 | 8.4 | 9.3 |
Hydrogen | 0.4 | 0.3 | 0.3 | 22.7 | 22.5 | 22.5 | 0.2 | 0 | 0 |
Subtotal | 80.6 | 80.2 | 79.9 | 148.3 | 143.8 | 142.4 | 109.6 | 108.1 | 107.0 |
Emissions Cost | 272.3 | 272.0 | 269.8 | 0 | 0 | 0 | 13.3 | 11.6 | 11.6 |
Total | 352.8 | 352.2 | 349.7 | 148.3 | 143.8 | 142.4 | 123.0 | 119.7 | 118.6 |

### v2

@@ 4 Results
In contrast, the zero-emissions scenarios have starkly greater variability, with 40-45 percent of delivered MWh having a zero or near-zero marginal cost ( $1/MWh). The large share of “free" energy occurs because solar and wind capacities are optimally built out to help serve demand on days with low to medium wind and sunlight and high demand. These times are seasonal, tending to occur in late fall and winter. As a result, there are often surpluses during other seasons, especially spring and summer, when wind and sunlight are more abundant, leading to many stretches of time with zero marginal cost. During these times, renewables are curtailed because supply exceeds demand and batteries are full. There can be zero-cost times without simultaneous curtailment if there is more energy in batteries than could be depleted before being fully charged again with excess wind and solar.
@@
In zero-emissions scenarios, the spread of marginal cost above near-zero is lumpy, with just 10 percent of MWh priced between $1 and $90, 20-30 percent spread evenly between $90 and $100 per MWh, 20-25 percent spread between $100 and $200 per MWh. Marginal costs above $200 are rare and extremely rare above $1000/MWh. Transmission mainly increases the frequency of hours in the near zero and $100 to $200 ranges and decreases them in the $90 to $100 range. Again, the influence of transmission is modest.
In the socially optimal scenarios, the distribution looks like a mixture of the least-cost and zero-emissions scenarios, with about 20 percent of MWh priced near zero, about 50 percent in the $20-50 range, and 25 percent in the $80 to $200 range.
@@
In all three cases, the influence of transmission on the distribution of marginal costs is modest. The extreme variability of marginal cost in higher-renewable scenarios, especially the high frequency of zero-marginal-cost energy, suggests that flexible demand could be valuable in ways quite unlike conventional systems where only critical peaks matter (imelda2022real).
@@
Transmission expansion increases economic rents for some resource owners while reducing them for others (Figure 6). If resource owners are paid a price equal to the the incremental cost of power in each region and hour, then all variable and fixed costs are recovered in long-run equilibrium, which is what the model assumes. In addition, some owners of scarce resources glean rent above their total costs. The potential rents we can identify in this study pertain to heterogeneous solar and wind resources and potentially constrained transmission resources. Areas with a comparative advantage in clean resources, such as unusually large or well-timed amounts of solar radiation or wind, earn rent because they provide more inframarginal value than marginal resources do. These rents are logically highest in the zero-emissions scenarios.
@@
Combining rent and marginal cost changes, which mostly benefit consumers, all regions gain from optimal transmission expansion. However, since some resource owners lose rents relative to the no-expansion case, diverging stakeholder interests could hamper efforts to expand transmission. Figure 6 illustrates these divergent interests. The graph shows gains and losses in solar, wind, nuclear, and transmission rent when going from the ZE scenario to the ZO scenario. Some rent gains are intuitive, such as wind in SPSS, the Rocky Mountain regions, and MISW, and rent losses in other, less windy areas. Gains in solar rents in NWPP and PJMW and loss of solar rents in CASO may be more surprising, but less so if one considers time patterns of net demand in these regions relative to neighbors. MISS loses rents for all three resources due to greater imports, with much greater benefits from transmission going to consumers (Figure 2). Nuclear plants generally lose rent from greater transmission, as it brings greater competition from solar and wind and much less is selected for expansion. (Rent accrues to existing nuclear if not expanded, as capital costs for existing are assumed sunk.) Existing transmission owners also lose substantial rents, as pointed out by hausman2024. The losses considered here, however, are forward-looking, and account for emerging substitutes and substantial projected demand growth.
It is interesting to compare rent implications of transmission expansion under full decarbonization versus social optimization, reported in the supplement (Figure A6). The geographic allocations are similar, but tend to be smaller in magnitude, especially for nuclear. Transmission rent losses are more concentrated in the Northeastern regions.
@@
Hourly operational detail in ERCOT shows how transmission expansion can aid resiliency during the most constrained times (Figure 7). The model optimizes hourly operations and interregional flows in each region, revealing how the system achieves balance. Here we consider costly day in a zero-emissions scenario in the most isolated region, TRE, the Eastern part of ERCOT, which has most of the interconnect’s generation and demand. In two of the four scenarios (LO, ZE), the day depicted (December 13th, 2050, based on actual weather data in 2012), is the most expensive in the region. This is a cold day in late Fall near the solstice, when solar resources would generally be more limited, and with unusually little wind.
@@
Compared with the least-cost scenarios, decarbonization is achieved via substantially greater capacities of solar and wind paired with significant use of batteries and hydrogen, which replaces natural gas and some old-vintage coal. The deficit from renewable energy is mostly counterbalanced with natural gas in the least-direct-cost scenarios and with imports, battery storage, and hydrogen in the zero-emissions scenarios. Notably, even the least-direct-cost scenarios use batteries to help serve the net peak at 7pm, when wind and solar power fall to near-zero.
@@
The effects of optimizing transmission are subtle and difficult to discern in the least-direct-cost scenarios. In these scenarios, some power is imported from the Western part of ERCOT (TREW) at 1 pm and 3 pm of December 13th, 2050 when transmission is fixed (panel a), but less is imported under optimal transmission, because, in this case, more natural gas capacity is built so the region can export more on other days. The generation mix is largely the same when transmission is fully optimized (panel b), except there is slightly more gas generation and less battery use during the evening peak. The peak hour marginal cost falls from about $15,000 to about $6,000 when optimizing transmission.
@@
The zero-emissions scenarios (c and d) show a somewhat greater influence from optimizing transmission. Generation mixes are similar regardless of transmission, but there is noticeably more battery use and slightly less wind in the scenario with existing transmission. And while imports help, especially in the optimal transmission scenario, it is interesting that imports occur midday to help replenish batteries, and much less during the expensive evening peak. The peak marginal cost of about $35,000 per MWh falls to about $20,000 with expanded transmission. And where the nighttime hours after the peak stay elevated at $5,000 per MWh with existing transmission, marginal cost falls back to very low levels when transmission is optimally expanded.
@@
Although we do not consider demand response or demand-side flexibility in this study, this difficult day in ERCOT suggests it could be valuable. In all scenarios, hours up to 5 pm are very low cost or free. Adjustments in EV charging, thermal storage or pre-heating in buildings, or even modest thermostat adjustments during the critical peak, would reduce required hydrogen or battery capacity, saving substantial costs. Even existing demand responses by some industrial customers could make a difference. Such demand responses would likely reduce transmission value, too (roberts2022).
@@ 5 Discussion
While the optimal expansion of interregional transmission would nearly quadruple what presently exists, we find that the net benefits of this expansion are modest relative to the overall cost of a decarbonized system. Alternatives to expanded transmission are multifaceted but mainly involve more battery storage, increased renewable generation capacity, and an expansion of nuclear power. While expanded transmission might improve resiliency in some locations, the overall distribution of marginal costs (i.e., competitive wholesale prices) is similar regardless of interregional transmission expansion. Consistent with wongel2023, our findings suggest many ways to firm solar and wind power, even using a more granular account of resources and an explicit account of transmission.
In an earlier draft of this paper, we found that optimal transmission expansion reduced the cost of decarbonization by 4% instead of 7%. We made a few revisions and updates, one of which made a discernible difference: we disallowed flexible ramping of nuclear power. Because nuclear is expensive but slightly cost reducing in decarbonized scenarios, its inflexibility makes transmission more valuable. In practice, nuclear power, especially newly built plants, can be more flexible than we assume. Partly for consistency and partly to err on the side of making transmission more valuable, we limit flexibility for all nuclear generation. For this reason and others indicated below, we likely overestimate the value transmission expansion.
@@
One qualification to these findings is that some kinds of transmission might be expanded at lower cost using various grid-enhancing strategies111https://www.energy.gov/oe/grid-enhancing-technologies-improve-existing-power-lines to improve power flow or by employing new high-capacity conductors in existing right-of-ways (chojkiewicz2024), which could halve the cost of expanding existing lines. It is also conceivable that storage costs could fall more than projected. Such technologies could only expand the options available to integrate clean energy affordably.
@@
Another qualification is that these findings assume economic efficiency or perfectly competitive outcomes. During congested times, transmission and generation capacity owners may possess considerable market power and be tempted to withhold generation or transmission to raise wholesale prices and profits (borenstein2002). Greater transmission capacity could, therefore, have the added benefit of increasing competition. At the same time, those earning substantial scarcity and/or market-power rents may resist expansion efforts that would reduce such rents. Transmission expansion may need to be expanded much more than otherwise optimal if the goal is to attenuate market power.
@@
Perhaps a greater limitation of this study is that it does not account for demand-side adjustments that could substitute for transmission, storage, and generation. Such responses could be achieved through strategic timing of electric vehicle charging, vehicle-to-grid technologies, improved building insulation, thermal storage, smart control of building HVAC and water heating systems, thermostat adjustments during extreme weather events, and industrial demand shifting, some of which already exist. While the demand data used in this study may already embody some demand responses from large commercial and industrial customers, the difficult-to-serve times in a high-renewable setting occur at very different times than in the baseline data. Thus, even existing demand response capabilities could reduce system costs and transmission relative to our estimates.
Given the extraordinary growth in marginal cost variability in the decarbonization scenarios, utilities, industry, communities, and opportunistic intermediaries should find ways to exploit growing wholesale price differentials. If they do, interregional transmission growth could be worth much less than we estimate. Future research could consider how optimal transmission upgrades and expansion interact with demand response and distributed resources from local to interregional scales. A robust demand-side market responsive to real-time marginal costs would also reduce market power during congested times.
At the end of 2023, more than 2.5 TW of solar, wind, and storage capacity was waiting for interconnection to the U.S. grid, a staggering quantity that, according to estimates here, could decarbonize the electricity sector projected for 2050. Still, the wait times are substantial, and many proposed projects are never completed (rand2023). Although this study does not focus on these immediate transmission challenges of grid interconnection and optimal power flow (which considers exact existing lines, their impedances, currents, voltage limits, and real and reactive power flows), it does account at least approximately for the cost of spur lines and transmission upgrades to connect new solar and wind installations to the grid. This study clarifies that while such local transmission upgrades are essential, these interconnection challenges should not be conflated with larger-scale interregional transmission upgrades. While substantial interregional transmission upgrades could be useful, there are also affordable substitutes.
@@ 6 Methods
Our analysis uses Switch (fripp2012switch; johnston2019switch), which is an open source capacity expansion model that minimizes the net present value (NPV) of costs for an electricity grid across all investment periods and timepoints. It optimizes both the investment and operational costs subject to constraints such as emissions limits or prices, or other policies, such as Renewable Energy Portfolio Standards (RPS).
@@
One of the most difficult parts of running electricity capacity expansion models is assembling all the data. PowerGenome (Powergenome) serves as a platform to generate inputs files for power system optimization models—including Switch, GenX, Temoa and USENSYS. The source data comes from a number of different resources, including EIA, NREL, and EPA.
@@
Existing generating units are from the latest version of form EIA-860 (EIA), 2021 supplemented with 860m from June 2023. The cost and heat rates of new-build resources are provided by NREL’s Annual Technology Baseline (ATB) 2022 (nrelatb2022). Renewable resources are generally location-specific, with unique generation profiles and interconnection costs. Rather than representing all potential new-build renewable resources as individual sites, PowerGenome lets the user specify how much capacity of each resource type should be available for consideration in a model region and how many clusters the resource should be represented by. We use 942 clusters of candidate renewable energy projects, corresponding to approximately 21 TW of potential capacity. High-resolution weather data provided by Vibrant Clean Energy for 2012 (the year upon which our demand data is based) is used to construct hourly variable capacity factors for each cluster.
@@
Existing transmission capacities between IPM regions are from EPA. Model regions in our study can consist of one or more individual IPM regions. When two or more IPM regions are combined to make a Switch model region, the transmission capacity between individual IPM regions is also combined. Transmission lines in this study are defined as connections between major metros in each region, with additional backbone networks connecting major metros within a region (if there is more than one). Costs and line loss use transmission line segments created with a least cost path method described in Appendix F of (patankar2023land).
@@
Hourly demand starts with NREL EFS profiles EFS. Stock values (historical) and hourly demand profiles of electrified end-use technologies like transportation, water heating, and space heating/cooling (derived from NREL EFS data) are subtracted from the 2019 EFS profiles. The remaining demand is inflated using sector-specific growth rates from EIA AEO 2022 (nalley2022annual). The future hourly demand from electrified end-use technologies are added back in using future stock values (from the REPEAT (REPEAT) scenario IRA_MID) and hourly demand profiles derived from EFS.
Contingency Demand
In high-renewable settings, peak demand times are not typically the most difficult to serve, but rather net peaks, which equal demand less supply from intermittent renewables, solar and wind. In systems with a lot of storage, even the highest net peak days may be relatively easy to serve if adjacent hours or days have lower net peaks. To test the model’s implications for transmission value in a robust way, we therefore found the highest three-day, demand-weighted average marginal cost in each region in the ZE scenario, and exogenously increased demand for all hours of these three days by 25%. Because transmission is heavily constrained in the ZE scenario, these three days tend to differ across regions, so boosting demand on these days ought to heavily favor transmission relative to baseline demand projections described above. We then resolve the zero-emissions scenarios, which we label BZE and BZO. Results for these scenarios are reported in the supplement (Figure A8). Within such system, the model predicts only a slightly higher value of optimal transmission expansion; it reduces overall costs by 8% instead of 7% relative to the existing transmission scenario.
Limitations
Three aspects are important to discuss in the context of transmission line expansion. First, we model transmission capacity in an aggregated manner to represent important energy flows between load zones. The existing capacities we employ are consistent with those in EPA’s Integrated Planning Model.222See https://www.epa.gov/power-sector-modeling/documentation-integrated-planning-model-ipm-base-case-v410. The regions in our model are aggregations of regions in the IPM model and transmission capacities between regions match aggregated capacities from their documentation. Our approach models the transfer capability of the electrical network but does not directly represent the physics of AC power flow. Hence, our findings should not be taken as a claim that there are few benefits associated with transmission expansion within individual regions in 2050. Our work sheds light on what are the most important electricity transfers between regions and interconnects to achieve zero emissions in 2050. Our findings speak to the modest economic impact of deploying “energy corridors” between interconnects rather than the value of specific transmission lines, for which power flow modeling would be required with an accurate model of the transmission network (e.g., California should be modeled with its approximately 9,000 buses). Second, our study, while high-resolution, considers only a single weather year and coincident demand. We leave for future work incorporating several weather years of data. Third, we do not account for reserves or for extreme weather events that do not occur in a typical year. Our exercise examining the robustness of our conclusions to artificially inflated demand during the most difficult-to-serve days in each region shows that the general conclusions may be robust to such considerations, but specific investments in each region would require a more granular look at the particular lines and more years of coincident weather and demand data to select the most prudent portfolio of options.
Figures
Optimal transmission expansion modestly reduces decarbonization costs of U.S. electricity
@@
Here we describe the mathematical structure of Switch in more detail and present additional results, with an emphasis on comparing the least-direct-cost (LX) scenarios with the socially optimal (SX) scenarios that minimize costs subject to a carbon emissions price of $190 per ton rather than forcing zero emissions. This price leads to 90 to 95% emissions reductions relative to 2022 levels in most cases, even with a 75% growth in electricity production. We refer to these scenarios as socially optimal because they equate the marginal cost of abatement to the Environmental Protection Agency’s marginal social cost of emissions (i.e., the estimated marginal benefit of abatement).
@@
Here we provide a brief overview of the mathematical formulation of Switch 2.0, the model used for our research. More complete documentation of the software can be found in johnston2019switch. Switch 2.0 has a modular architecture that reflects the modularity of actual power systems. Most power system operators follow rules that maintain an adequate supply of power, and most individual devices are not concerned with the operation of other devices. Similarly, core modules in Switch define spatially and temporally resolved balancing constraints for energy and reserves, and an overall social cost. Separate modules represent components such as generators, batteries or transmission links. These modules interact with the overall optimization model by adding terms to the shared energy and reserve balances and the overall cost expression. Switch 2.0 supports co-optimization of multiple investment periods, but we have omitted those definitions here, since we use a single stage for this study (2041-2050). We have likewise omitted details on spinning reserves and unit commitment, which were not used for this study.
@@
Additional terms define behavior of storage facilities (charging, state of charge, round-trip losses), fuel consumption (full load heat rate times power production) and hydrogen facilities (production and storage of hydrogen and conversion back to electricity). For complete details on Switch’s mathematical formulation, see the Supplementary Material of johnston2019switch at https://ars.els-cdn.com/content/image/1-s2.0-S2352711018301547-mmc1.pdf.
@@
Least-direct-cost | Zero-emissions | Socially-optimal | Supplemental | |||||||||
LE | LI | LO | ZE | ZI | ZO | SE | SI | SO | BE | BO | ZE+ | |
Panel A: Cost components (Bil $) | ||||||||||||
Fuels | 37.6 | 36.8 | 36.5 | 6.4 | 4.4 | 4.2 | 23.7 | 22.1 | 21.4 | 6.1 | 3.7 | 5.6 |
OM | 4.1 | 4.0 | 4.0 | 2.0 | 1.4 | 1.4 | 6.9 | 6.8 | 6.5 | 1.9 | 1.2 | 1.8 |
Hydrogen | 0.9 | 0.8 | 0.7 | 28.6 | 29.0 | 28.2 | 0.7 | 0.4 | 0.4 | 35.6 | 30.6 | 27.9 |
Storages | 1.0 | 0.8 | 0.8 | 16.6 | 13.0 | 12.5 | 4.4 | 3.5 | 3.4 | 17.9 | 13.4 | 16.9 |
Generations | 50.0 | 49.4 | 49.0 | 125.0 | 112.4 | 110.6 | 84.7 | 83.4 | 82.8 | 125.5 | 109.8 | 120.1 |
Transmissions | 3.2 | 4.5 | 4.7 | 3.2 | 12.8 | 13.9 | 3.2 | 8.0 | 8.7 | 3.2 | 15.9 | 5.2 |
SUBTOTAL | 96.8 | 96.3 | 95.9 | 181.9 | 173.0 | 170.7 | 123.6 | 124.1 | 123.2 | 190.2 | 174.7 | 177.5 |
Emissions | 332.3 | 327.7 | 326.8 | 0.0 | 0.0 | 0.0 | 14.6 | 9.8 | 9.9 | 0.0 | 0.0 | 0.0 |
TOTAL | 429.1 | 424.0 | 422.7 | 181.9 | 173.0 | 170.7 | 138.2 | 133.9 | 133.1 | 190.2 | 174.7 | 177.5 |
Panel B: Marginal cost ($/MWh) | ||||||||||||
Load-weighted-MC | 37.9 | 37.6 | 37.6 | 70.4 | 66.1 | 65.5 | 54.8 | 51.5 | 51.1 | 71.7 | 66.2 | 69.3 |